OpenStudy (anonymous):

how many integral values of n are possible for (\sqrt{n}+\sqrt{n+2013}) to be rational?

OpenStudy (anonymous):

it is \[\sqrt{n}+\sqrt{n+2013}\] !!

OpenStudy (anonymous):

n could be anything greater than zero I think...

OpenStudy (anonymous):

what values ??

OpenStudy (aravindg):

for a number to be rational we should be able to express it in form a/b where a and b are 2 integers , clearly \[x \ge 0 \] for square root to operate

OpenStudy (anonymous):

Yes, what Aravind said.

OpenStudy (aravindg):

its *n sorry

OpenStudy (aravindg):

now both \(\sqrt{n}\) and \(\sqrt{n+2}\) are irrational numbers ...but we can get a rational numbers for some particular values of n clearly n not equal to 0 now see we which is the next perfect square after 2013 ?

OpenStudy (aravindg):

*n+2013 sorry for typo

OpenStudy (anonymous):

next perfect square after 2013 is \[45^2\]=2025?

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

u der?

OpenStudy (aravindg):

pls wait a min

OpenStudy (anonymous):

ok!

OpenStudy (aravindg):

i was trying to get to the solution using trial and error method ...finding a suitable square number that satisfies our situation

OpenStudy (anonymous):

so what would be those numbers?

OpenStudy (anonymous):

by hit and trial n can be \[196\] and what else?

OpenStudy (anonymous):

also it would be very long !

OpenStudy (aravindg):

i understand ... i am trying an alternate solution ...will post if I succeed ..thanks for patience I got guests at my house sorry for delay